# Luminous Fast Blue Optical Transients as very massive star core-collapse events

### v1

@@ Acknowledgements.
AAC acknowledges support through the European Space Agency (ESA) research fellowship programme. P.G.J. is supported by the European Union (ERC, Starstruck, 101095973, PI Jonker). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them.This work made use of v2.2.1 of the Binary Population and Spectral Synthesis (bpass) models as described in Eldridge et al. (2017) and Stanway & Eldridge (2018). This work has made use of ipython (Perez & Granger, 2007), numpy (Harris et al., 2020), scipy (Virtanen et al., 2020); matplotlib (Hunter, 2007), Seaborn packages (Waskom, 2021) and astropy,(https://www.astropy.org) a community-developed core Python package for Astronomy (Astropy Collaboration et al., 2013; Price-Whelan et al., 2018).

### v2

@@ Acknowledgements.
We thank the referee Brian Metzger for their insightful comments on this manuscript. AAC acknowledges support through the European Space Agency (ESA) research fellowship programme. P.G.J. is supported by the European Union (ERC, Starstruck, 101095973, PI Jonker). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council Executive Agency. Neither the European Union nor the granting authority can be held responsible for them.This work made use of v2.2.1 of the Binary Population and Spectral Synthesis (bpass) models as described in 2017PASA...34...58E and 2018MNRAS.479...75S. This work has made use of ipython (2007CSE.....9c..21P), numpy (2020Natur.585..357H), scipy (2020NatMe..17..261V); matplotlib (2007CSE.....9...90H), Seaborn packages (Waskom2021) and astropy,(https://www.astropy.org) a community-developed core Python package for Astronomy (astropy:2013; astropy:2018).
Appendix A Additional figures
In this Appendix we provide two additional figures. Figure 7 shows the progenitor temperatures, luminosities and radii for the fiducial selection of BH progenitors where the BH mass exceeds 38 M⊙. Figure 8 shows the carbon-oxygen core versus BH masses for the selected models.
Appendix B Disc radii
To calculate the initial disc radii - for input to the model of 2020MNRAS.492.5655M, 2025MNRAS.541..429M - we first define M⋆=MMleftover. The initial disc radius can be estimated from angular momentum and mass conservation as the star is “turned” (from some process we do not specify) into a Keplerian flow. We therefore need an estimate of the maximum angular momentum available to the disc. On energetics grounds we can write for the pre-collapse star,
where is the maximum pre-collapse stellar rotational velocity. The corresponding specific angular momentum of the star is approximately . Substituting into equation B.1 (and ignoring constants of order unity) we find
and therefore
The angular momentum in the accretion disc has to be less than or equal to the total pre-collapse angular momentum, therefore
Where and is the initial disc radius. We therefore find an absolute maximum333In principle the initial disc radius could be larger than this if the disc material somehow got all of the initial stellar angular momentum but not all of the leftover mass. We ignore this possibility here. initial disc radius of
in the extreme case that (i) the star is maximally rotating and (ii) the total pre-collapse angular momentum is transferred entirely to the disc. Since and we can introduce factors to allow for sub-critical rotation velocities , and the fraction of angular momentum which goes into the disc , as follows:
Finally, we adopt for the pre-collapse rotational velocity and introduce the fraction of leftover mass which is accreted , yielding
as adopted in Section 4.3.
The disc then must expand, as angular momentum is conserved while the mass of the flow drops (as some material is added to the black hole). To be precise, the quantity
is conserved. The mass in the disc drops with time in a self similar fashion, as the accretion rate
at large times (e.g. Pringle81), such that
and thus
Of course, this expansion of the disc can be suppressed if angular momentum is lost from the disc – i.e., if angular momentum is transported to large radii by winds. It is important therefore that the radii at large times (in this model) are larger than that inferred from observations.